OpenStudy (anonymous):

sinx cscx -cscx+2sinx=0

OpenStudy (apoorvk):

sinx cscx = 1 therefore, 1- (1/sinx)+ 2sinx=0 so, [2(sinx)^2 + sinx -1]sinx =0. we can eliminate sinx in denominator but as sinx can be zero, and then equation will then become infinity. so basically we will write a pre-note that x is unequal to any n(pi), where n is an integer. next in the the equation we substitute sinx with any variable 't' and solve like any quadratic, then the solutions of sinx will come 2t^2 +t -1 = 0 we get t= sinx = 1/2 or -1. so x can be pi/6 or -pi/2 in th eprinciple values, or 2n(pi) added to the prinicipal values for a general set of solutions.

OpenStudy (anonymous):

thank you. can please help me with this problem sin 3x cos x - cos 3x sin x=.5

OpenStudy (apoorvk):

okay easy. (sin a) (cos b) + (cos a) (sin b) = sin(a+b) so here th eaboce equation gets down to sin(3x+x)= 0.5 sin inverse of half = 4x so 4x = 30 degrees so x = 7.5 degrees.

OpenStudy (anonymous):

tanx cotx -tanx + 2cotx= 0

OpenStudy (apoorvk):

try this one now its similar to the first one that you asked!! in fact exactly the same procedure!

OpenStudy (anonymous):

sin 3x cos x - cos 3x sin x=.5 you gave me the answer to positive cos 3x not negative

OpenStudy (apoorvk):

oops sorry. my bad. then the formula condenses to sin(A-B) as you may have gusesed so, sin(3x-x) = sin2x = 0.5 2x= 30 degrees so x= 15 degrees.